Convert (a) 50 rev to radians, (b) 48 rad to revolutions, (c) 72 rps(revolutions per second) to degree/s, (d ) 1.50 x 103 rpm (revolution per minute) to rad/s,


1
Expert's answer
2021-06-22T11:31:04-0400

(a) "50\ rev=50\cdot2\pi=314.159\ rad"


(b) "48\ rad=7.639 \ rev"


(c) "72\ rps=72\cdot360°=25920\ degree/s"


(d) "1.50\cdot10^3\ rpm=1500\cdot2\pi/60=157.08\ rad/s"